rahul runs 420 metres in 48 seconds .what distance will he cover in 12 seconds.
Answer by dkppathak(439)

48 second he will go 420 mtrs
in 1 second it will go=420/48=105/12
in 12 second he willl go =12x105/12
answer 105 meters